OpenStudy (anonymous):

Given the linear equation y = -4x – 2, find the y-coordinates of the points (1, ), (-1, ), and (-3, ). Please show all of your work. Plot these points and graph the linear equation

OpenStudy (anonymous):

Just plug in those numbers for x. FOr example, I'll do the first one. y = -4x - 2 y = -4(1) - 2 y = -4 - 2 y = -6 (1, -6) See? Do the same for the others.

OpenStudy (anonymous):

So the next one would be y=-4(-1)-2 y=4-2 y=2 (1, 2)

OpenStudy (anonymous):

Yes, except that the final answer for that would be (-1, 2) since the x-value changed.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

And the last one I ended up with (10, -3)

OpenStudy (anonymous):

Close. It would be (-3, 10). You got the values mixed up. I assume you know how to plot these and just draw a line through them?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

Can you help me with this one? Find the x- and y-intercepts for the equation -x + 2y = 4

OpenStudy (anonymous):

Set x = 0 to get the y-intercept and set y = 0 to get the x-intercept.

OpenStudy (anonymous):

So -0 + 2(0)=4?

OpenStudy (anonymous):

Set them equal to zero separately.

OpenStudy (anonymous):

-(0) + 2y = 4 -x + 2(0) = 4 Just solve for the x and y in each of those. Y-Intercept: (0, ?) X-intercept: (?, 0)

OpenStudy (anonymous):

So y =2?

OpenStudy (anonymous):

Yup. y-intercept: (0, 2)

OpenStudy (anonymous):

So x intercept would be (2, 0) ?

OpenStudy (anonymous):

No solve: -x + 2(0) = 4 Then put the answer where the question mark is: (?, 0)

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

Simplifying it further: -x + 2(0) = 4 -x + 0 = 4 -x = 4 x = -4

OpenStudy (anonymous):

Ohh okay thank you

OpenStudy (anonymous):

No Problem :)